OpenStudy (anonymous):

1/log8 (12)-log8 (3) evaluate the expression

OpenStudy (imer):

\[\frac{ 1 }{ [\log_{8}(12)-\log_{8}(3)] } =\frac{ 1 }{ [\log_{8}(4*3)-\log_{8}(3)] }=\frac{ 1 }{ [(\log_{8}(4)+\log_{8}(3))-\log_{8}(3)] }\] \[\frac{ 1 }{ [\log_{8}(4)] }\] But we know \[8^{x}=4\] and we need to find "x" \[(2^3)^{x}=2^2\] Therefore, "x" has to be "2/3" to cancel 3. \[\frac{ 1 }{\frac{ 2 }{ 3 } }=\frac{ 3 }{ 2 }\]